# Jailbreaking Multimodal Large Language Models via Shuffle Inconsistency

### v1

@@ Appendix A Hyper-parameter of Patch Nums
Table A.1: Attack performance for different patch nums in SI-Attack. The results are based on the sub-dataset (01-Illegal-Activity) in MM-safetybench (without typograhpy).
@@ Appendix B Influence of Text Shuffling Types
Here we explore different ways of text random shuffling operations including: no shuffling, shuffling all the words, shuffling only nouns and adj, shuffling trigrams, and shuffling within trigrams. And the corresponding results can be viewed in Table B.1. We can see that randomly shuffling all the words can obtain the best results, so we select this shuffling approach in our final attack setting.
Table B.1: Attack performance for different text shuffling types in SI-Attack. The results are based on the sub-dataset (01-Illegal-Activity) in MM-safetybench (without typograhpy).
@@
All 3.96 80.41
@@ Appendix C Performance on Different Scales’ MLLMs
Table C.1: Attack performance for different scales’ MLLMs in SI-Attack. We select different scales’ versions of the InternVL-2, including 4B, 8B, and 26B. The results are based on the sub-dataset (01-Illegal-Activity) in MM-safetybench (without typograhpy).
@@
Appendix D More Results on MM-safetybench
Here we conduct the MM-safetybench without harmful typography, and the results can be viewed in Table D.1 and Table D.2. From the results, our SI-Attack can obviously enhance the attack effectiveness compared with the QR Attack for both the open-source and closed-source models. Specifically, for the open-source models of LLaVA-NEXT, MiniGPT-4, InternVL-2, and VLGuard, our SI-Attack achieves attack success rates of 37.98%, 54.88%, 48.15%, and 39.88%, which are better than the original jailbreak attack instructions 19.77%, 33.81%, 34.82%, and 25.49%, respectively; for the closed-source models of GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Qwen-VL-Max, our SI-Attack can increase the attack success rate by 35.95%, 32.21%, 32.14%, and 38.69%, respectively.
Table D.1: Results of Query-Relevant Attack (QR) and our SI-Attack in the metric of toxic score (Toxic) and attack success rate (ASR%) on open-source MLLMs. The harmful instructions are based on MM-safetybench (without typography) and evaluated by ChatGPT-3.5. “01-IA” to “13-GD” denote the 13 sub-dataset of prohibited scenarios, and the “ALL” denotes the results on the whole harmful instructions.
@@ Appendix D More Results on MM-safetybench
Attack QR-Attack[15] SI-Attack QR-Attack[15] SI-Attack QR-Attack[15] SI-Attack QR-Attack[15] SI-Attack
@@
Table D.2: Results of Query-Relevant Attack (QR) and our SI-Attack in the metric of toxic score (Toxic) and attack success rate (ASR%) on closed-source MLLMs. The harmful instructions are based on MM-safetybench (without typography) and evaluated by ChatGPT-3.5. “01-IA” to “13-GD” denote the 13 sub-dataset of prohibited scenarios, and the “ALL” denotes the results on the whole harmful instructions.
@@
Attack QR-Attack[15] SI-Attack QR-Attack[15] SI-Attack QR-Attack[15] SI-Attack QR-Attack[15] SI-Attack
@@
Appendix E Harmful Instruction Case of SI-Attack
Here we provide different types of harmful cases, including the generated harmful images without typography, generated harmful images with typography, and images with only harmful typography, which can be found in Figure E.1, Figure E.2, and Figure E.3, respectively.
Figure E.1: The Claude-3.5-Sonnet responses for the origin harmful instructions and harmful instructions optimized by SI-Attack. And the harmful images are combined without harmful typography. The harmful instruction pairs are based on MM-safetybench.
Figure E.2: The Claude-3.5-Sonnet responses for the origin harmful instructions and harmful instructions optimized by SI-Attack. And the harmful images are combined with harmful typography. The harmful instruction pairs are based on dataset of HADES.
Figure E.3: The Claude-3.5-Sonnet responses for the origin harmful instructions and harmful instructions optimized by SI-Attack. And the harmful images only contain harmful typography. The harmful instruction pairs are based on SafeBench.
Appendix F Designed Prompt for Toxic Score Judge
Following [26], we provide the automatic identification prompt for toxic score evaluation of the MLLMs’ responses in Figure F.1, which mainly contains the prohibited activities of OpenAI and the toxic score criteria. We need to provide the origin question and MLLM’s response for the prompt and query the toxic judge model ChatGPT-3.5.
Figure F.1: The Automatic identification prompt for toxic score evaluation designed for judge ChatGPT-3.5.

### v2

@@ Appendix A Hyper-parameter of Patch Nums
Table A.1: Attack performance for different patch nums in SI-Attack. The results are based on the sub-dataset (01-Illegal-Activity) in $M$-safetybench (without typograhpy).
@@ Appendix B Influence of Text Shuffling Types
Here we explore different ways of text random shuffling operations including: no shuffling, shuffling all the words, shuffling only nouns and adj, shuffling trigrams, and shuffling within trigrams. Meanwhile, we also explore the token-wise shuffle based on the BPE tokenizer, which is wisely applied in GPT series models. And the corresponding results can be viewed in Table B.1. We can see that randomly shuffling all the words can obtain the best results, so we select this shuffling approach in our final attack setting.
Table B.1: Attack performance for different text shuffling types in SI-Attack. The results are based on the sub-dataset (01-Illegal-Activity) in $M$-safetybench (without typograhpy).
@@
Token-wise Shuffle 3.95 72.16
Word-wise Shuffle 3.96 80.41
@@ Appendix C Performance on Different Scales’ MLLMs
Table C.1: Attack performance for different scales’ MLLMs in SI-Attack. We select different scales’ versions of the InternVL-2, including 4B, 8B, and 26B. The results are based on the sub-dataset (01-Illegal-Activity) in $M$-safetybench (without typograhpy).
@@
Appendix D Adaptive SI-Attack against PPL Detector
There is a type of method [alon2023detecting, jain2023baseline] that detects the text perplexity and then judges whether the text has attack intention. Here we apply an adaptive attack method for this type of defense method. We first perform perplexity detector on the shuffled harmful texts before attack optimization. Only when the perplexity detector is passed will the shuffled harmful texts starts the attack optimization process. To make it easier for texts to pass the perplexity detection, we adopt a Trigram-based text shuffling operation, while the other settings remain the same as the original settings. Here we apply the Llama-3.1 [dubey2024llama] as the perplexity detector instead of GPT-2, while other settings keep the same with [alon2023detecting]. The experiments in Table D.1 show that in the face of perplexity detection defense, our method still maintains a competitive attack performance, which shows the generalization and scalability of our SI-Attack.
Table D.1: Adaptive SI-Attack performance against PPL detector. The results are based on the sub-dataset (01-Illegal-Activity) in $M$-safetybench (without typograhpy).
Attack Target Model Toxic Score ASR(%)
Baseline GPT-4o 2.51 35.05
GPT-4o+PPL Detector 2.51 35.05
SI-Attack GPT-4o 3.96 80.41
GPT-4o+PPL Detector 3.83 71.13
Appendix E More Results on MM-safetybench
Here we conduct the MM-safetybench without harmful typography, and the results can be viewed in Table E.1 and Table E.2. From the results, our SI-Attack can obviously enhance the attack effectiveness compared with the QR Attack for both the open-source and closed-source models. Specifically, for the open-source models of LLaVA-NEXT, MiniGPT-4, InternVL-2, and VLGuard, our SI-Attack achieves attack success rates of 37.98%, 54.88%, 48.15%, and 39.88%, which are better than the original jailbreak attack instructions 19.77%, 33.81%, 34.82%, and 25.49%, respectively; for the closed-source models of GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Qwen-VL-Max, our SI-Attack can increase the attack success rate by 35.95%, 32.21%, 32.14%, and 38.69%, respectively.
Table E.1: Results of Query-Relevant Attack (QR) and our SI-Attack in the metric of toxic score (Toxic) and attack success rate (ASR%) on open-source MLLMs. The harmful instructions are based on MM-safetybench (without typography) and evaluated by ChatGPT-3.5. “01-IA” to “13-GD” denote the 13 sub-dataset of prohibited scenarios, and the “ALL” denotes the results on the whole harmful instructions.
@@ Appendix E More Results on MM-safetybench
Attack QR-Attack[liu2023mm] SI-Attack QR-Attack[liu2023mm] SI-Attack QR-Attack[liu2023mm] SI-Attack QR-Attack[liu2023mm] SI-Attack
@@
Table E.2: Results of Query-Relevant Attack (QR) and our SI-Attack in the metric of toxic score (Toxic) and attack success rate (ASR%) on closed-source MLLMs. The harmful instructions are based on MM-safetybench (without typography) and evaluated by ChatGPT-3.5. “01-IA” to “13-GD” denote the 13 sub-dataset of prohibited scenarios, and the “ALL” denotes the results on the whole harmful instructions.
@@
Attack QR-Attack[liu2023mm] SI-Attack QR-Attack[liu2023mm] SI-Attack QR-Attack[liu2023mm] SI-Attack QR-Attack[liu2023mm] SI-Attack
@@
Appendix F Harmful Instruction Case of SI-Attack
Here we provide different types of harmful cases, including the generated harmful images without typography, generated harmful images with typography, and images with only harmful typography, which can be found in Figure F.1, Figure F.2, and Figure F.3, respectively.
Figure F.1: The Claude-3.5-Sonnet responses for the origin harmful instructions and harmful instructions optimized by SI-Attack. And the harmful images are combined without harmful typography. The harmful instruction pairs are based on $M$-safetybench.
Figure F.2: The Claude-3.5-Sonnet responses for the origin harmful instructions and harmful instructions optimized by SI-Attack. And the harmful images are combined with harmful typography. The harmful instruction pairs are based on dataset of HADES.
Figure F.3: The Claude-3.5-Sonnet responses for the origin harmful instructions and harmful instructions optimized by SI-Attack. And the harmful images only contain harmful typography. The harmful instruction pairs are based on SafeBench.
Appendix G Designed Prompt for Toxic Score Judge
Following [wang2024multi-jail], we provide the automatic identification prompt for toxic score evaluation of the MLLMs’ responses in Figure G.1, which mainly contains the prohibited activities of OpenAI and the toxic score criteria. We need to provide the origin question and MLLM’s response for the prompt and query the toxic judge model ChatGPT-3.5. This evaluation is conducted to simultaneously verify the effectiveness and harmfulness of the model’s responses. Only when the model’s responses are toxic and fully respond to the attacker’s intent, can the highest score be obtained, which can be applied to judge both the comprehension ability and safety ability for MLLMs.
Figure G.1: The Automatic identification prompt for toxic score evaluation designed for judge ChatGPT-3.5.